The monthly profit from the sale of a product is given by P = 18x − 0.1x2 − 100 dollars.

(a) What level of production maximizes profit?
(b) What is the maximum possible profit?
Answer #1

Solution :-

Given data :

The monthly profit from the sale of a product is given by P = 18x − 0.1x2 − 100 dollars.

(a) What level of production maximizes profit?

Here consider differentiation of P with respect to x, we get

dPd da dx 102-100) (18x 0.12

dar- r(0.1x2)--Ir (100) (18 r)

dP ー=18-0.1*21

dP 18-0.2.r

Again consider differentiation with respect to x, we get

rac{d^2P}{dx^2}=rac{d}{dx}left {rac{dP}{dx} ight }

rac{d^2P}{dx^2}=rac{d}{dx}left {18-0.2x ight }

18)0.2)

P =-12 < 0

Here,P will be maximum at critical point. i.e  rac{dP}{dx}=0

Rightarrow 18 - 0.2 x =0

Rightarrow 18 = 0.2 x

Rightarrow  18 0.2

Rightarrow x = 90

RightarrowTherefore, level of production maximizes profit ,x = 90.

(b) What is the maximum possible profit?

Per 182-0.12.2-100

Paar = 18(90) _ 0.1(90)2-100

P, 1620-810-100

Pmar- 710

RightarrowThe maximum possible profit, Pmar- 710